OpenStudy (anonymous):

Given the 1 in. width of the hex nut, what is the length of one of the six sides?

OpenStudy (anonymous):

@jim_thompson5910 what do i need to d o?

jimthompson5910 (jim_thompson5910):

What kind of width is it? Width Across Flats? or Width Across Corners? (see attached picture to get an idea of what I'm asking)

OpenStudy (anonymous):



OpenStudy (anonymous):

width across corners

jimthompson5910 (jim_thompson5910):

Sorry the drawing feature doesn't work on this site for some reason

jimthompson5910 (jim_thompson5910):

hint: see attached image

OpenStudy (anonymous):

the drawing is showing the face of the nut with a line going up the side of it on the right saying 1 in. so i would say that is across corners maybe

OpenStudy (anonymous):

the first thing i would do is take 90-0.5 correct?

jimthompson5910 (jim_thompson5910):

no, you would need to look at a 30-60-90 triangle (since 90 - 60 = 30)

jimthompson5910 (jim_thompson5910):

here's a template of one http://upload.wikimedia.org/wikipedia/commons/4/45/30-60-90_triangle.jpg

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

1/2*0.5=0.25

jimthompson5910 (jim_thompson5910):

Good, that's the value of x. There are 2 x values that line each side, so each side is 2x = 2(0.25) = 0.5 units long

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

np